OpenStudy (anonymous):

1) Prove that the base angles of an isosceles triangle are acute angles. 2) Find the measures of an isosceles triangle if the base angle is twice bigger than the vertex angle.

OpenStudy (rulnick):

isosceles implies two base angles are identical call their measures x sum of base angles is 2x sum of all three angles is 180 call third angle's measure y 2x + y = 180 2x = 180 - y but y>0 so 2x < 180 x < 90

OpenStudy (anonymous):

1. all three angles in a triangle is 180. in a isoleces the base angles equal each other. i agree with rulnick

OpenStudy (rulnick):

For #2, x+2x+2x = 180 5x = 180 x= 36 36, 72, 72 is the ans (degrees)

OpenStudy (anonymous):

how about if the base angle is three times smaller than its exterior angle

OpenStudy (rulnick):

then the base angle would have to be 90, which is impossible for isosceles

OpenStudy (anonymous):

ok, thank you so much